# Bridging the Modality Gap: Enhancing Channel Prediction with Semantically Aligned LLMs and Knowledge Distillation

### v1

@@ Abstract
Accurate channel prediction is essential in massive multiple-input multiple-output (m-MIMO) systems to improve precoding effectiveness and reduce the overhead of channel state information (CSI) feedback. However, existing methods often suffer from accumulated prediction errors and poor generalization to dynamic wireless environments, making it challenging to maintain high prediction accuracy. Large language models (LLMs) have demonstrated remarkable modeling and generalization capabilities in tasks such as time series prediction, making them a promising solution. Nevertheless, a significant modality gap exists between the linguistic knowledge embedded in pretrained LLMs and the intrinsic characteristics of CSI, posing substantial challenges for their direct application to channel prediction. Moreover, the large parameter size of LLMs hinders their practical deployment in real-world communication systems with stringent latency constraints. To address these challenges, we propose a novel channel prediction framework based on semantically aligned large models, referred to as CSI-ALM, which bridges the modality gap between natural language and channel information. Specifically, we design a cross-modal fusion module that aligns CSI representations with the language feature space using a pretrained corpus. Additionally, we maximize the cosine similarity between word embeddings and CSI embeddings to construct semantic cues, effectively leveraging the latent knowledge in LLMs. To reduce complexity and enable practical implementation, we further introduce a lightweight version of the proposed approach, called CSI-ALM-Light. This variant is derived via a knowledge distillation strategy based on attention matrices, which extracts essential features from the teacher model, CSI-ALM, and transfers them to a compact, efficient student model, CSI-ALM-Light. Extensive experimental results demonstrate that CSI-ALM achieves a 1 dB gain over state-of-the-art deep learning methods. Moreover, under limited training data conditions—where all models are trained using only 10% of the original training dataset—CSI-ALM-Light, with only 0.34M parameters, attains performance comparable to CSI-ALM and significantly outperforms conventional deep learning approaches. These validate the effectiveness of the proposed approach for accurate and efficient channel prediction in m-MIMO systems.
@@ I Introduction
The recent rise of Transformer-based architectures and large language models (LLMs) [30] has created new opportunities for channel prediction in wireless communications. Although originally designed for natural language tasks, LLMs possess a powerful ability to model long-range dependencies and sequential data, making them well-suited for capturing the temporal dynamics of wireless channels with improved accuracy and generalization. This cross-domain potential has led to several promising applications. For example, [31] introduces an LLM-based beam prediction method using textual prompts, achieving superior robustness and performance compared to traditional LSTM models. In [32], LLMs are used for resource allocation, where fine-tuning on a small dataset delivers results comparable to advanced reinforcement learning techniques. Additionally, [33] explores a distributed training framework combining federated learning and aerial computing to support the deployment of large models in communication systems.
Recent studies have begun exploring the application of pre-trained large language models (LLMs) in channel prediction. For example, [37] proposes an LLM-based method for downlink channel prediction, where specialized embedding layers are designed for the frequency and angular domains while keeping the LLM backbone fixed. In [38], the downlink CSI prediction task is reformulated to align with the autoregressive “next-token” generation mechanism of LLMs, resulting in a CSI-LLM framework that supports variable-length inputs and multi-step continuous prediction. However, these approaches do not fully address the fundamental differences between linguistic and channel data modalities. Directly applying pre-trained LLMs to channel prediction tasks may underutilize their representational capabilities, limiting the effective exploitation of the embedded knowledge within the models. Furthermore, the large parameter sizes and high computational complexity of LLMs present challenges for practical deployment in wireless communication systems.
@@
  • Extensive experiments show that CSI-ALM delivers a 1 dB performance improvement over the state-of-the-art methods. Furthermore, in scenarios with limited training data—using just 10% of the original dataset—CSI-ALM-Light, with only 0.34 million parameters, achieves accuracy comparable to CSI-ALM while substantially outperforming other baseline approaches. These results highlight the proposed framework’s effectiveness in enabling both accurate and efficient channel prediction for m-MIMO systems.
The rest of this article is organised as follows. Section II introduces the MIMO channel prediction model and performance metrics studied. Section III provides a detailed description of the proposed CSI-ALM framework and the modality alignment method. Section IV introduces the proposed CSI-ALM-Light and the self-attention relationship-based knowledge distillation method. Section V discusses the simulation results, and Section VI concludes the paper.
@@ II-A Channel Model
[…]
@@ II-B Downlink CSI Prediction
[…]
@@ III-A LLM based Pipeline
The input uplink CSI sequence first undergoes a series of preprocessing steps. However, directly feeding the high-dimensional sequence into the network can lead to substantial computational complexity and prolonged training times, particularly when the number of antennas and subcarriers is large. To address this efficiently, we propose a parallel antenna processing approach, where the CSI for each transmitter–receiver antenna pair $(i,j)$is predicted separately. Here, $i$ and $j$ denote the indices of the transmitter and receiver antennas, respectively. Specifically, the input sequence at time $t$ for each pair $\bm{\mathit{H}}^{t,T}_{u,ij}\in\mathbb{C}^{F\times T}$ is the CSI sequence between the $i$-th transmitting antenna and the $j$-th receiving antenna over the past $T$ time steps:
@@
where $\bm{\mathit{H}}^{t-p}_{u,ij}$ is the CSI of the transmitter antenna $i$ and the receiver $j$ at time $t-p$, $F$ denotes the number of subcarriers per antenna pair. We recall that $T$ is the temporal length of historical CSI used for prediction.
@@ IV-A Knowledge Distillation
The core of our method lies in leveraging the multi-head attention relations derived from the scaled dot-product of queries, keys, and values across multiple attention heads in both the teacher and student models for knowledge distillation. As illustrated in Fig.4, we first concatenate the queries, keys, and values from different attention heads, and then split the concatenated vector into individual components according to the desired number of relation heads. The same procedure is applied to the keys and values. This approach enables a more flexible handling of the self-attention mechanism, particularly in scenarios where the teacher and student models employ differing numbers of attention heads. After reorganizing the queries, keys, and values, we proceed to compute the attention relations of both models:
@@
where $A_{r}$ is the number of relation heads. If the number of relation heads and attention heads is the same, the Q-K relation is equivalent to the attention weights in the self-attention module. $\alpha_{ij}\in\{0,1\}$ is the weight assigned to each self-attention relation loss. We transfer query-query, key-key and value-value self-attention relations to balance the performance and training cost.
@@ IV-B CSI-ALM-Light
In order to meet the requirements of inference speed and model size in practical deployment, we need to design small yet sophisticated student model, called CSI-ALM-Light. As shown in Fig.6, for the design of the CSI-ALM-Light, we also adopt the Transformer architecture. Compared to the complex structure of the CSI-ALM, only a simple CSI embedding module and the final projection layer are retained here. Similar to CSI-ALM, we also use a parallel approach to process antennas here, inputting $\bm{\mathit{H}}^{t,T}_{u,ij}$passes through the embedding layer to obtain CSI embedding:
@@ VI-A Experimental Setup
We have considered the 3GPP Urban Macro (UMa) channel model. The initial position of the user is randomized. We set the training set, validation set, and test set to 9000, 1000, and 1000 samples respectively, with user speeds evenly distributed between 10km/h and 100km/h. For the training and validation sets, there are 10 samples per speed point, while for the test set, there are 100 samples per speed point. For TDD and FDD systems, the center frequency of the uplink carrier is set to 2.4GHz. For FDD systems, we choose the carrier data adjacent to the uplink channel as the carrier frequency of the FDD system. The hidden layer dimension of the transformer layer in our student model is 128, and 3 transformer layers are stacked. The length of the soft prompt is set to 4.
We selected advanced neural network models for channel prediction, including CNN, RNN, LSTM, GRU, and Transformer, as our baseline methods. In the experiment, we directly used the CSI values of the uplink channel over the next L steps as the predicted CSI values for the downlink channel. For the GPT-2 model in our proposed method, we opted for the smallest GPT-2 variant, with a hidden layer dimension of 768 in the transformer layers, and deployed the first six layers. The length of semantic prompts is set to 4.
@@ VI-B CSI Prediction Performance
Additionally, considering TDD system, we evaluate the zero-shot Compared to the 3GPP Uma channel scenario, the Umi scenario has more complex environmental characteristics. The Umi scenario has a smaller coverage area and higher user density, with base stations deployed at lower heights. Additionally, the channel fading and multipath effects in the Umi scenario are more pronounced, and issues such as co-channel interference and neighboring cell interference are more prominent. Therefore, we evaluate the zero-shot generalization capability of all models by training them in the Uma scenario and directly applying them to the Umi scenario, while keeping all other settings unchanged. The results in Fig.13 show that, during model transfer testing in the TDD system, the LLM-based method retains the strong generalization capability observed in NLP tasks. Furthermore, our CSI-ALM framework, equipped with modality alignment, further improves cross-scenario prediction performance, outperforming other methods. We then conducted the same transfer testing in the FDD system. As shown in Fig.14, due to the strong frequency selectivity of the channel, all methods exhibit poor NMSE performance. However, our proposed approach still achieves the smallest NMSE, significantly outperforming other methods. This is because our proposed modality alignment scheme can better leverage the pre-trained knowledge of LLM and align the features between textual knowledge and CSI data with a small amount of training data.
Figure 15: Results of ablation experiment.
TABLE I: Inference speed and complexity comparison.
@@ VI-C Ablation Experiment
The fundamental advantage of LLM-based CSI prediction over traditional methods lies in its ability to leverage pretrained knowledge. To further analyze this aspect, we compare three different initialization strategies: (1) random initialization, (2) initialization from a pretrained base model, and (3) removal of the LLM backbone. We conducted ablation experiments under TDD system, as illustrated in Fig. 15, initializing with a pretrained base model yields the best performance, significantly outperforming the other two approaches. The randomly initialized model lacks access to pretrained knowledge, rendering the modality alignment mechanism ineffective. Consequently, its performance becomes comparable to that of the model without the LLM backbone. These results further confirm the critical role of pretrained knowledge in enhancing the effectiveness of channel prediction.
@@ VI-D Model Lightweighting Experiment
In this section, we conduct experiments on the deployment of lightweight models. Specifically, we use a knowledge distillation method based on self attention relationships to obtain CSI-ALM-Light, whose parameters are similar to those of a small deep learning model. We train the model using 10% of the original training set and evaluate the predictive performance of CSI-ALM-Light and other small deep learning models under TDD and FDD systems. As shown in Fig. 16 and 17, when the amount of training data is small, although our CSI-ALM-Light has a simple structure, its NMSE performance in predicting CSI is still better than other traditional deep learning methods. In addition, we observe a very interesting phenomenon: when only 10% of the training data is used, the performance of our student model was almost the same as that of the teacher model. This is because, in the case of limited data, CSI-ALM often cannot learn the general patterns of the data, while CSI-ALM-Light, guided by well-trained CSI-ALM, can learn richer prior knowledge and store it in the attention layer. By refining the self attention relationship, the CSI-ALM-Light can quickly learn the prior knowledge of the CSI-ALM, thereby achieving better performance. At the same time, we present the performance of training the student model separately without knowledge distillation. It can be seen that the predictive performance of training the student model separately is far inferior to that of the CSI-ALM-Light that has undergone knowledge distillation, whether in TDD or FDD scenarios. This further demonstrates the effectiveness of our proposed knowledge distillation method based on self attention relationships.
@@
VI-E Training and Inference Cost
We also compare the parameter counts and inference times of different models to evaluate their deployment challenges in practical scenarios. Our experiment was conducted on a single RTXA6000. As shown in Table I, it can be observed that although LLM-based approaches inherently exhibit a larger parameter count due to their architectural characteristics, their inference speed remains competitive with smaller deep learning models owing to specialized acceleration mechanisms. Furthermore, the distilled student model addresses the challenge of excessive parameter size and achieves comprehensive optimization in both parameter efficiency and inference latency.

### v2

@@ Abstract
Accurate channel prediction is essential in massive multiple-input multiple-output (m-MIMO) systems to improve precoding effectiveness and reduce the overhead of channel state information (CSI) feedback. However, existing methods often suffer from accumulated prediction errors and poor generalization to dynamic wireless environments, making it challenging to maintain high prediction accuracy. Large language models (LLMs) have demonstrated remarkable modeling and generalization capabilities in tasks such as time series prediction, making them a promising solution. Nevertheless, a significant modality gap exists between the linguistic knowledge embedded in pretrained LLMs and the intrinsic characteristics of CSI, posing substantial challenges for their direct application to channel prediction. Moreover, the large parameter size of LLMs hinders their practical deployment in real-world communication systems with stringent latency constraints. To address these challenges, we propose a novel channel prediction framework based on semantically aligned large models, referred to as CSI-ALM, which bridges the modality gap between natural language and channel information. Specifically, we design a cross-modal fusion module that aligns CSI representations with the language feature space using a pretrained corpus. Additionally, we maximize the cosine similarity between word embeddings and CSI embeddings to construct semantic cues, effectively leveraging the latent knowledge in LLMs. To reduce complexity and enable practical implementation, we further introduce a lightweight version of the proposed approach, called CSI-ALM-Light. This variant is derived via a knowledge distillation strategy based on attention matrices, which extracts essential features from the teacher model, CSI-ALM, and transfers them to a compact, efficient student model, CSI-ALM-Light. Extensive experimental results demonstrate that CSI-ALM consistently outperforms state-of-the-art deep learning methods across various communication scenarios, achieving substantial performance gains. Moreover, under limited training data conditions—where all models are trained using only 10% of the original training dataset—CSI-ALM-Light, with only 0.34M parameters, attains performance comparable to CSI-ALM and significantly outperforms conventional deep learning approaches. These validate the effectiveness of the proposed approach for accurate and efficient channel prediction in m-MIMO systems.
@@ I Introduction
The recent rise of Transformer-based architectures and large language models (LLMs) [30] has created new opportunities for channel prediction in wireless communications. Although originally designed for natural language tasks, LLMs possess a powerful ability to model long-range dependencies and sequential data, making them well-suited for capturing the temporal dynamics of wireless channels with improved accuracy and generalization. This cross-domain potential has led to several promising applications. For example, [31] introduces an LLM-based beam prediction method using textual prompts, achieving superior robustness and performance compared to traditional LSTM models. In [32], LLMs are used for resource allocation, where fine-tuning on a small dataset delivers results comparable to advanced reinforcement learning techniques. Additionally, [33] explores a distributed training framework combining federated learning and aerial computing to support the deployment of large models in communication systems. [34] and [35] investigate the construction of efficient semantic communication systems using multimodal large models. Unlike conventional distortion-free transmission schemes, their approaches leverage large models to achieve high semantic fidelity while transmitting only minimal information. [36] further proposes a framework that integrates LLMs with multi-agent systems, enabling knowledge retrieval, task decomposition, and feedback mechanisms in 6G communications, thereby enhancing the intelligence and adaptability of communication systems.
Recent studies have begun exploring the application of pre-trained LLMs in channel prediction. For example, [37] proposes an LLM-based method for downlink channel prediction, where specialized embedding layers are designed for the frequency and angular domains while keeping the LLM backbone fixed. In [38], the downlink CSI prediction task is reformulated to align with the autoregressive “next-token” generation mechanism of LLMs, resulting in a CSI-LLM framework that supports variable-length inputs and multi-step continuous prediction. However, these approaches do not fully address the fundamental differences between linguistic and channel data modalities. Directly applying pre-trained LLMs to channel prediction tasks may underutilize their representational capabilities, limiting the effective exploitation of the embedded knowledge within the models. Furthermore, the large parameter sizes and high computational complexity of LLMs present challenges for practical deployment in wireless communication systems.
@@
  • Extensive experimental results demonstrate that CSI-ALM consistently outperforms state-of-the-art deep learning methods across various communication scenarios, achieving substantial performance gains. Furthermore, in scenarios with limited training data—using just 10% of the original dataset—CSI-ALM-Light, with only 0.34 million parameters, achieves accuracy comparable to CSI-ALM while substantially outperforming other baseline approaches. These results highlight the proposed framework’s effectiveness in enabling both accurate and efficient channel prediction for m-MIMO systems.
The rest of this article is organised as follows. Section II introduces the MIMO channel prediction model and performance metrics studied. Section III provides a detailed description of the proposed CSI-ALM framework and the modality alignment method. Section IV introduces the proposed CSI-ALM-Light and the self-attention relationship-based knowledge distillation method. Section V introduces the model training method. Section VI discusses the simulation results, and Section VII concludes the paper.
@@ II-A Channel Model
$\bm{\mathit{H}}(f)=\int_{\theta\in\mathcal{A}_{k}}\alpha(\theta)\mathbf{a}(\theta)e^{-j2\pi f\tau(\theta)+j\phi(\theta)}d\theta,$ (1)
@@ II-B Downlink CSI Prediction
$\text{s.t.}\bm{\hat{\mathit{H}}}_{d}^{t}=g({\bm{\mathit{H}}_{u}^{t-1}},{\bm{\mathit{H}}_{u}^{t-2}},\dots,{\bm{\mathit{H}}_{u}^{t-T}}),$ (5)
@@ III-A LLM based Pipeline
The input uplink CSI sequence first undergoes a series of preprocessing steps. However, directly feeding the high-dimensional sequence into the network can lead to substantial computational complexity and prolonged training times, particularly when the number of antennas and subcarriers is large. To address this efficiently, we propose a parallel antenna processing approach, where the CSI for each transmitter–receiver antenna pair (i,j)(i,j)is predicted separately. Here, $i$ and $j$ denote the indices of the transmitter and receiver antennas, respectively. Specifically, the input sequence at time $t$ for each pair $\bm{\mathit{H}}^{t,T}_{u,ij}\in\mathbb{C}^{F\times T}$ is the CSI sequence between the $i$-th transmitting antenna and the $j$-th receiving antenna over the past $T$ time steps:
@@
where $\bm{\mathit{H}}^{t-p}_{u,ij}$ is the CSI of the transmitter antenna $i$ and the receiver $j$ at time tpt-p, $F$ denotes the number of subcarriers per antenna pair. We recall that $T$ is the temporal length of historical CSI used for prediction.
@@ IV-A Knowledge Distillation
The core of our method lies in leveraging multi-head attention relations, derived from the scaled dot-product of queries, keys, and values across multiple attention heads, for knowledge distillation between the teacher and student models. As illustrated in Fig.4, we first concatenate the queries, keys, and values from different attention heads, and then split the concatenated vector into individual components according to the desired number of attention heads. The same procedure is applied to the keys and values. This approach enables a more flexible handling of the self-attention mechanism, particularly in scenarios where the teacher and student models employ differing numbers of attention heads. After reorganizing the queries, keys, and values, we proceed to compute the attention relations of both models:
@@
where $A_{r}$ is the number of relation heads. If the number of relation heads and attention heads is the same, the Q-K relation is equivalent to the attention weights in the self-attention module. $\alpha_{xy}\in\{0,1\}$ is the weight assigned to each self-attention relation loss. We transfer query-query, key-key and value-value self-attention relations to balance the performance and training cost.
@@ IV-B CSI-ALM-Light
In order to meet the requirements of inference speed and computational complexity in practical deployment, we introduce a small yet sophisticated student model, called CSI-ALM-Light. As shown in Fig.6, CSI-ALM-Light adopts the same decoder-only architecture as GPT-2. Compared with the more complex structure of CSI-ALM, CSI-ALM-Light uses fewer hidden state dimensions and layers, retaining only a simple CSI embedding module and the final projection layer. Similar to CSI-ALM, we also use a parallel approach to process antennas here, inputting $\bm{\mathit{H}}^{t,T}_{u,ij}$passes through the embedding layer to obtain CSI embedding:
@@ VI-A Experimental Setup
We consider the 3GPP Urban Macro (UMa) channel model, with the initial user position randomized. The training, validation, and test sets contain 9000, 1000, and 1000 samples respectively, with user speeds uniformly distributed between 10km/h and 100km/h. In both the training and validation datasets, 10 distinct samples are generated for each velocity, while in the test dataset, each velocity is associated with 100 independent samples to ensure a more comprehensive evaluation of model generalization. For both TDD and FDD systems, the center frequency of the uplink carrier is set to 2.4GHz. In FDD systems, the carrier frequency is chosen as the band adjacent to the uplink channel. During training, we employ the Adam optimizer with a batch size of 1024 and a learning rate of $1.0\mathrm{e}{-4}$ . During training, we employ the Adam optimizer with a batch size of 1024 and a learning rate of $1.0\mathrm{e}{-4}$.
We select advanced neural network models for channel prediction, including CNN, RNN, LSTM, GRU, and Transformer, as our baseline methods. The hidden layer dimension of the RNN, LSTM, and GRU is set to 256, with a total of 3 layers. The CNN model consists of 10 convolutional layers with a kernel size of 3×3. The Transformer model uses an encoder-decoder structure with a hidden layer dimension of 384, 3 encoder layers, and 2 decoder layers. In the experiment, we directly used the CSI values of the uplink channel over the next L steps as the predicted CSI values for the downlink channel. Since our goal is to demonstrate the effectiveness of our proposed cross-modal semantic alignment method, and to balance performance and computational cost for experimental purposes, we select the smallest GPT-2 variant as the LLM backbone network in our proposed method. Its Transformer layer has a hidden dimension of 768 and the first six layers are deployed. The length of the semantic hint is set to 4. Performance improves with increasing model parameters, and our method is also applicable to other more advanced pre-trained LLMs, such as GPT-3 and LLaMA-7B. For CSI-ALM-Light, the architecture adopts a Decoder-only Transformer consisting of 3 hidden layers, each with a hidden size of 128 and 8 attention heads. The length of the soft prompt is set to 4.
@@ VI-B CSI Prediction Performance
TABLE I: Average NMSE performance of each method in TDD and FDD scenarios.
Scenario CSI-ALM LLM4CP Transformer CNN GRU LSTM RNN
TDD -13.62dB -12.61dB -11.32dB -11.81dB -10.27dB -10.59dB -10.27dB
FDD -3.51dB -3.09dB -2.25dB -0.24dB -2.73dB -2.32dB -2.48dB
@@
Figure 15: The zero-shot generalization performance for the FDD systems in highway scenario.
Additionally, considering TDD system, we evaluate the zero-shot Compared to the 3GPP Uma channel scenario, the Umi scenario has more complex environmental characteristics. The Umi scenario has a smaller coverage area and higher user density, with base stations deployed at lower heights. Additionally, the channel fading and multipath effects in the Umi scenario are more pronounced, and issues such as co-channel interference and neighboring cell interference are more prominent. Therefore, we evaluate the zero-shot generalization capability of all models by training them in the Uma scenario and directly applying them to the Umi scenario, while keeping all other settings unchanged. The results in Fig.13 show that, during model transfer testing in the TDD system, the LLM-based method retains the strong generalization capability observed in NLP tasks. Furthermore, our CSI-ALM framework, equipped with modality alignment, further improves cross-scenario prediction performance, outperforming other methods. We then conducted the same transfer testing in the FDD system. As shown in Fig.14, due to the strong frequency selectivity of the channel, all methods exhibit poor NMSE performance. However, our proposed approach still achieves the smallest NMSE, significantly outperforming other methods. This is because our proposed modality alignment scheme can better leverage the pre-trained knowledge of LLM and align the features between textual knowledge and CSI data with a small amount of training data. To more clearly demonstrate the numerical results, we also present the average NSME values of each model under both TDD and FDD scenarios in Table I. As shown in Table I, in TDD scenarios the proposed CSI-ALM method achieves an average gain of over 1 dB compared with the state-of-the-art LLM4CP method. In the FDD scenario, despite the increased channel complexity and the greater difficulty of CSI feature extraction, the proposed method still achieves an average performance gain of about 0.4 dB, demonstrating its robustness under challenging conditions.
Finally, we further evaluate the generalization capability of the proposed method under high-mobility communication scenarios. In such FDD environments, the pronounced Doppler effect often causes conventional algorithms to suffer from a rapid performance degradation. As illustrated in the Fig. 15, our proposed method demonstrates stronger robustness in this setting, offering an effective solution for channel prediction in extreme communication environments.
TABLE II: Inference speed and complexity comparison.
@@ VI-C Ablation Experiment
In this section, we conduct extensive ablation experiments to verify the effectiveness of each module in our CSI-ALM. The fundamental advantage of LLM-based CSI prediction over traditional methods lies in its ability to leverage pretrained knowledge. To further analyze this aspect, we compare three different initialization strategies: (1) random initialization, (2) initialization from a pretrained base model, and (3) removal of the LLM backbone. We conducted ablation experiments under TDD system, as illustrated in Fig. 16, initializing with a pretrained base model yields the best performance, significantly outperforming the other two approaches. The randomly initialized model lacks access to pretrained knowledge, rendering the modality alignment mechanism ineffective. Consequently, its performance becomes comparable to that of the model without the LLM backbone. These results further confirm the critical role of pretrained knowledge in enhancing the effectiveness of channel prediction.
Figure 16: Results of ablation experiment.
Moreover, as shown in Fig. 16, emoving the cross-modal attention module and semantic prompts, both designed to align CSI features with textual knowledge, leads to a significant degradation in channel prediction performance. Among these two factors, the impact of removing the cross-modal attention module is particularly pronounced. In fact, without this module, performance drops to a level comparable to that of randomly initialized LLM backbones. This observation further demonstrates that, without effective modality alignment, the pretrained knowledge embedded in LLMs cannot be fully leveraged, making it difficult to enhance channel prediction performance.
TABLE III: NMSE performance of CSI-ALM-light with different model sizes.
NMSE Parameter inference time
-11.19dB 0.349M 2.41ms
-11.42dB 0.449M 2.46ms
-11.64dB 0.548M 2.52ms
-11.91dB 0.648M 2.84ms
@@ VI-D Model Lightweighting Experiment
In this section, we conduct experiments on the deployment of lightweight models. Specifically, we use a knowledge distillation method based on self attention relationships to obtain CSI-ALM-Light, whose parameters are similar to those of a small deep learning model. We train the model using 10% of the original training set and evaluate the predictive performance of CSI-ALM-Light and other small deep learning models under TDD and FDD systems. As shown in Fig. 17 and 18, When the amount of training data is small, in most speed scenarios, although our CSI-ALM-Light has a simple structure, its NMSE performance in predicting CSI is still better than other traditional deep learning methods. In addition, we observe a very interesting phenomenon: when only 10% of the training data is used, the performance of our student model was almost the same as that of the teacher model. This is because, in the case of limited data, CSI-ALM often cannot learn the general patterns of the data, while CSI-ALM-Light, guided by well-trained CSI-ALM, can learn richer prior knowledge and store it in the attention layer. By refining the self attention relationship, the CSI-ALM-Light can quickly learn the prior knowledge of the CSI-ALM, thereby achieving better performance. At the same time, we present the performance of training the student model separately without knowledge distillation. It can be seen that the predictive performance of training the student model separately is far inferior to that of the CSI-ALM-Light that has undergone knowledge distillation, whether in TDD or FDD scenarios. This further demonstrates the effectiveness of our proposed knowledge distillation method based on self attention relationships.
At the same time, although CSI-ALM-Light performs worse than baseline methods under certain speed conditions in TDD scenarios, this is mainly due to model capacity limitations. We also evaluate CSI-ALM-Light of model different sizes in TDD scenarios. As shown in the Table II, even a small increase in model capacity leads to a significant improvement in performance. This implies the model can be flexibly scaled according to system requirements to balance deployment efficiency and prediction accuracy.
@@
VI-E Modal Alignment Validity Analysis
In this section, we provide a thorough analysis of the effectiveness of our proposed modality alignment method. As mentioned earlier, the pre-trained LLM inherently operates within the linguistic semantic space, as it is trained entirely on a text corpus. Therefore, when directly applied to physical domain representations such as CSI, it fails to effectively leverage its rich semantic priors, resulting in a significant modality gap. To address this issue, we introduce a modality alignment strategy based on a cross-modal attention mechanism, enabling the model to project CSI features into a shared latent semantic space compatible with the LLM pre-trained linguistic representations. To further validate the effectiveness of this mechanism, we visualize the feature distributions of CSI embeddings and pre-trained word embeddings for LLM4CP and our proposed CSI-ALM, as shown in Fig. 19 and 20. The results clearly reveal fundamental differences in representational behavior. Specifically, the CSI embeddings generated by LLM4CP tend to form multiple isolated clusters that are significantly separated from the word embeddings. This scattered distribution indicates that LLM4CP fails to establish effective alignment between CSI embeddings and the semantic space of the pre-trained language model, resulting in suboptimal knowledge transfer and, in turn, degraded prediction performance. In contrast, after applying the proposed modality alignment strategy, the CSI embeddings generated by CSI-ALM have a distribution closer to the pre-trained word embeddings and exhibit stronger structural cohesion. The emergence of these compact and semantically related clusters indicates that the CSI representations are effectively aligned with the pre-trained language feature space. This improved cohesion enables the LLM backbone to interpret CSI features in a manner consistent with its learned language structure, thereby enhancing semantic understanding and facilitating cross-domain knowledge transfer. Furthermore, the increased density of aligned embeddings reflects a reduction in modality discrepancy, implying that the model learns a more unified and semantically meaningful latent representation space. By bridging the gap between physical-domain CSI data and language semantics, our approach enables the pre-trained LLM to leverage its prior knowledge for channel prediction tasks, thereby improving the model’s interpretability, robustness, and generalization across diverse communication environments.
Figure 19: t-SNE visualization of LLM4CP pre-trained word Embeddings and CSI Embeddings.
Figure 20: t-SNE visualization of CSI-ALM pre-trained word Embeddings and CSI Embeddings.
VI-F Analysis of the Effectiveness of Knowledge Distillation
In this section, we provide a more detailed analysis to further demonstrate the effectiveness of the proposed knowledge distillation strategy. Specifically, we visualize the attention heatmaps from the last layer of CSI-ALM, CSI-ALM-light, and CSI-ALM-light without knowledge distillation. As shown in Fig. 21 and 22, the attention patterns of CSI-ALM and CSI-ALM-light exhibit a high degree of similarity, indicating that the CSI-ALM-light successfully inherits the structural knowledge captured by the CSI-ALM. This resemblance suggests that the distillation process effectively preserves the key contextual dependencies learned by the teacher model, enabling the lightweight model to maintain comparable interpretability while significantly reducing computational cost. In constrast, as shown in Fig. 23, the attention patterns of the model trained without knowledge distillation appear disordered and unstructured. These results demonstrate that the proposed distillation framework not only compresses the model but also faithfully preserves critical attention mechanisms, thereby validating the practical applicability of CSI-ALM-light in real-world channel prediction tasks.
VI-G Training and Inference Cost
We also compare the parameter counts and inference times of different models to evaluate their deployment challenges in practical scenarios. Our experiment was conducted on a single RTXA6000. As shown in Table II, it can be observed that although LLM-based approaches inherently exhibit a larger parameter count due to their architectural characteristics, their inference speed remains competitive with smaller deep learning models owing to specialized acceleration mechanisms. Furthermore, the distilled student model addresses the challenge of excessive parameter size and achieves comprehensive optimization in both parameter efficiency and inference latency.
Figure 21: Attention heatmap of the CSI-ALM.
Figure 22: Attention heatmap of the CSI-ALM-light.
Figure 23: Attention heatmap of the CSI-ALM-light without distillation.